5. In a materials testing laboratory, a metal wire made from a new alloy is found to break when a tensile force of 90.8 N is applied perpendicular to each end. If the diameter of the wire is 1.84 mm, what is the breaking stress of the alloy?

1
Expert's answer
2022-06-06T08:55:02-0400

"p=\frac FS=\frac{4F}{\pi d^2}=34~MPa."